The speed of sound v in a medium depends on its wavelength λ, the young modulus E, and the density ρ, of the medium. Use the method of dimensional analysis to derive a formula for the speed of sound in a medium
1
Expert's answer
2020-03-02T09:07:48-0500

"v=\lambda ^\alpha E^\beta \rho ^\gamma"


"[\lambda]=[m]=L"


"[E]=[Pa]=ML^{-1}T^{-2}"


"[\rho]=[kg/m^3]=ML^{-3}"


"[v]=[m\/s]=LT^{-1}"


"LT^{-1}=L^\alpha(ML^{-1}T^{-2})^\beta(ML^{-3})^\gamma"


We have


"\alpha-\beta-3\gamma=1"

"\beta+\gamma=0"

"-2\beta=-1"


The solution of the system gives


"\alpha=0"

"\beta=1/2"

"\gamma =-1/2"


"v=\lambda ^0 E^{1/2} \rho ^{-1/2}=\sqrt{\frac{E}{\rho}}" Answer.